OpenStudy (anonymous):

The data in the table relates the total cost y to have a party with x guests at a restaurant.   Number Of Guests              Total Cost($)                  X                                      Y                 10                                     200                 15                                     300                 25                                   500                 30                                   600 Based on the information in the table, what is the function rule? y = x + 190 x = y + 190 y = 20x x = 20y

OpenStudy (anonymous):

What do you do to x to make y?

OpenStudy (anonymous):

It is y=20x because all you do is look at it and tell that x affects the outcome of y so it would be y= Then you see which ones are true. Like y=20x you would test it by going through the numbers and seeing if it fits. y=20*10 is 100. y=20*15 is 300 y=20*25 is 500 and y=20*30 is 600

OpenStudy (anonymous):

Oh ok. Thank for the explanation too it helps so that I can use it for help on othe questions!:D Thank You!

OpenStudy (anonymous):

Still working on this one????

OpenStudy (anonymous):

Na I finished with this one! :D

OpenStudy (zzr0ck3r):

why does it keep getting bumped?

OpenStudy (anonymous):

I think I might have done it by accident haha oops :z :D

OpenStudy (zzr0ck3r):

you can close it:)

OpenStudy (anonymous):

Oh ok I thought I did:D I'll check it right now :D

OpenStudy (zzr0ck3r):

no worries, was just not sure you knew.

OpenStudy (anonymous):

Oh yes I see it says open I just closed it : 0 :D thanks :D